A hectare is a unit of area equal to 10,000 square meters.
A 150-hectare plot of land is rectangular and is 1.2 kilometers in width. Give the perimeter of this land.

Answer by addingup(3677)

1km = 1000 meters
1.2*1000 = 1200 meters wide
:
The problem says: 1 hectare = 10,000m^2, and says there are 150 hectares:
150*10,000 = 1,500,000 square meters
:
The perimeter:
2L+2W = 1,500,000
We know W, so substitute:
2L+2(1200) = 1,500,000
2L+2,400 = 1,500,000
2L = 1497600
L = 748800 meters is the perimeter. If you want it in kilometers divide by 1,000
:
John